OpenStudy (anonymous):

Jupiter has a mass that is approximately 318 times greater than the earth. The acceleration due to gravity on Jupiter's surface is 2.64 times larger than that on Earth. From these data, obtain the radius of Jupiter expressed as a multiple of the earth's radius.

OpenStudy (mos1635):

M1=318*Mo g1=2.64*go g=G*M/R^2

OpenStudy (anonymous):

basically change the equation F=G*Mm/r^2. But what happens to the second mass ?

OpenStudy (mos1635):

g=F/m

OpenStudy (mos1635):

g1=G*M1/R1^2 go=G*Mo/Ro^2 divide.......

OpenStudy (anonymous):

Alright. Thank you.

OpenStudy (mos1635):

you are welcome